evening shift summary for 7/1/99

AG (ag@fnal.gov)
Fri, 2 Jul 1999 00:59:46 -0400 (EDT)

Shift Summary for eve shift, 7/1/99

On Shift: Peter Shanahan (PSh), Sasha Golossanov (AG)

Hours of beam delivered at all intensities: ~7.
Hours of beam delivered at nominal intensity: ~6.5
Hours of E832 data (special runs not included): ~7.
Hours of special runs (laser scans, muon runs...): .5

Accounting of lost time:

1. Accelerator down
2. Swapped DYC in crate 53
3. Booster vacuum tripping
4. Muon run

========================================================================

Runs ended this shift:

Run 13782 - 0.65M events - BEAM_832
Run 13783 - 58.8M events - BEAM_832
Run 13785 - 3.78M events - BEAM_832

Tapes brought to FCC this shift: none

========================================================================

Accesses during shift: none

========================================================================

Overview of shift:

More or less smooth 832 data-taking, starting ~16:00, with some minor
problems and fluctuations in the beam intensity, the latter probably
being the most bothersome problem for this shift.

========================================================================

Description of Shift:

run spill time comment

------------------------------------------------------------------------

13782 56-81 1611-1645 Persisting "PIPELINE" ERRORS => quick reset

Probably these errors are due to the DYC in
crate 53 swapped earlier.
We decide to put the old one back:
1. stopped run 13782
2. put the old DYC back
3. initialize new run.

13783 7 1705 No "PIPELINE" ERRORS anymore.

13783 60 1825 Load KQNL06 on ktev2

problem on ktev2 - process "15" takes
~70% more cpu time. this seems to be pid 23869,
which has accumulated 148 min of cpu time,
compared to ~85min for other
ktev2 l3_driver_filts.
( see ktev logbook XX, p.105 for details)

13783 2058 "Booster vacuum trip. Investigating"

stop run to:
1. re-initialize L3
2. try to get at least 1 pass laser scan to
investigate channel 961.

2105 Beam returns -> re-initialize BEAM_832

13785 2113 Start run 13785

2115 Beam goes away without explanation

2123 Beam returns

2130 Call MCR complain about intensity
fluctuations between ~ 5 and 10 sec
into spill.

2157 "Resetting extraction devices.
Estimate 10 minutes"

13785 40 2210 Beam is back - the structure is better.

13785 60 2240 seems like this process "15" problem is gone

13785 68 2246 intensity fluctuations are back again.

13785 102 2330 stop this run in order to make muon run.

13786 2350 KQs failed to initialize -> fresh start.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

CsI 961 Seems to be OK, but it should be watched closely.